# Mitigating Perception Bias: A Training-Free Approach to Enhance LMM for Image Quality Assessment

### v1

@@ Abstract
Despite the impressive performance of large multimodal models (LMMs) in high-level visual tasks, their capacity for image quality assessment (IQA) remains limited. One main reason is that LMMs are primarily trained for high-level tasks (e.g., image captioning), emphasizing unified image semantics extraction under varied quality. Such semantic-aware yet quality-insensitive perception bias inevitably leads to a heavy reliance on image semantics when those LMMs are forced for quality rating. In this paper, instead of retraining or tuning an LMM costly, we propose a training-free debiasing framework, in which the image quality prediction is rectified by mitigating the bias caused by image semantics. Specifically, we first explore several semantic-preserving distortions that can significantly degrade image quality while maintaining identifiable semantics. By applying these specific distortions to the query/test images, we ensure that the degraded images are recognized as poor quality while their semantics remain. During quality inference, both a query image and its corresponding degraded version are fed to the LMM along with a prompt indicating that the query image quality should be inferred under the condition that the degraded one is deemed poor quality. This prior condition effectively aligns the LMM’s quality perception, as all degraded images are consistently rated as poor quality, regardless of their semantic difference. Finally, the quality scores of the query image inferred under different prior conditions (degraded versions) are aggregated using a conditional probability model. Extensive experiments on various IQA datasets show that our debiasing framework could consistently enhance the LMM performance and the code will be publicly available.
1 Introduction
@@
Figure 1: Illustration of perception bias in Large Multimodal Model (LMM) during quality assessment. Image quality ratings from the LMM (mPLUG-Owl3 [56]) were obtained using the Q-Bench testing framework [49]. The LMM consistently assigns higher quality ratings to images in the second row compared to the first, despite both sets exhibiting similar quality distributions as measured by Mean Opinion Scores (MOSs). This discrepancy suggests that the LMM relies more on image semantics than on low-level image clues for quality assessment. Zoom in for enhanced clarity.
@@
Image Quality Assessment (IQA) models aim to measure image quality in alignment with human perception, playing a fundamental role across various computer vision tasks [48]. According to the availability of reference information, the IQA models can be classified into Full-Reference IQA (FR-IQA), Reduced-Reference IQA (RR-IQA), and No-Reference IQA (NR-IQA) [60]. Among these, NR-IQA presents particular values as the image quality can be predicted without a reference requirement. In the past decades, significant progress has been witnessed in NR-IQA, including traditional hand-crafted feature-based models [34, 16, 32, 31] and deep learning-based models [20, 64, 66, 13, 42]. Despite the advancement, current models still grapple with limited generalization capabilities on unseen scenes and distortions due to the significant distribution shifts between training and test sets. Recently, the emergence of Large Multimodal Models (LMMs) [37] has demonstrated impressive generalization abilities across various vision-language tasks, such as classification [47], image captioning [59], and visual question answering [40]. However, the focus on high-level visual tasks usually limits their effectiveness in low-level tasks, such as IQA [49, 43]. Although pioneering researchers have attempted to fine-tune or retrain LMMs for improved IQA accuracy [50, 8, 67], the laborious dataset construction and costly model training usually render this approach inefficient. Moreover, tuning LMMs specifically for IQA also introduces the risk of catastrophic forgetting [30], compromising the retention of general knowledge and ultimately degrading the models’ capability on other tasks.
@@
A powerful strategy to both retain the LMM’s strengths across tasks and enhance its IQA performance is to unlock its vast general knowledge through well-crafted prompts, encouraging the LMM to respond accurately to quality rating requests. However, despite this potential, LMMs—driven by training objectives that emphasize semantic extraction over quality—usually default to interpreting image quality heavily relying on image semantics. As shown in Fig. 1, two sets of distorted images with different semantics are presented to an LMM for quality rating. The results indicate that the LMM consistently prefers the quality of images with the second set of semantics, despite both sets having similar quality (Mean Opinion Score (MOS)) distributions. The case reveals that the LMM intrinsically bases its quality rating on image semantics rather than on quality-related clues (e.g., noise or blur).
@@
Motivated by this observation, we introduce an innovative, training-free approach to mitigate the perception bias inherent in LMMs. Our enhancement strategy consists of two main steps: 1) bias exposure, and 2) bias mitigation. In the bias exposure step, we assume that the same bias exists consistently across images sharing the same semantics. Based on this assumption, we can expose the perception bias of a query/test image by measuring how much the LMM resists labeling it as high quality when its quality is severely degraded but the semantics are preserved. To achieve this, we explore several specific distortions that drastically corrupt the image quality while preserving its semantics to some extent. For a query image, we then impose these distortions on the query image and obtain its degraded versions. Herein, those degraded images should be deemed as poor quality. However, the LMM may not always agree with the fact and the disagreement leads to the bias exposure. In the bias mitigation step, we address the exposed bias using an instructive conditional prompt. Specifically, during inference, we provide both the query image and its degraded version to the LMM, along with a prompt indicating that the quality of the query image should be assessed under the condition that the degraded counterpart is rated as poor quality. By forcing the LMM to rate the quality of the degraded images appropriately, the bias mitigation in turn refines the LMM’s quality prediction for the query image. Finally, the quality predictions under different distortions are aggregated through a conditional probability model, further improving the prediction accuracy. Before delving into detail, we highlight our main contributions as follows:
@@ 2 Related work
NR-IQA. Traditional NR-IQA methods mainly rely on handcrafted features to model natural scene statistics (NSS) and estimate image quality by assessing the destruction level of the NSS. For example, Mean-Subtracted Contrast-Normalized (MSCN) coefficients were explored to model NSS as a unified distribution, enabling quality estimation through distributional discrepancy analysis [34]. Such NSS descriptors were also widely investigated in the wavelet domain [35, 44], gradient domain [54] as well as the discrete cosine transform (DCT) domain [39]. Drawing inspiration from the free-energy brain theory that the human visual system (HVS) strives to minimize uncertainty through an internal generative model [12, 11], psychovisual quality models were proposed in [16, 17, 61]. Tough valuable insights from these studies, the inherent complexity of HVS continues to challenge feature design, prompting a shift toward data-driven feature learning. Kang et al. proposed the first convolutional neural networks (CNNs) for IQA and extended this work into a multi-task learning framework [20, 21]. Inspired by the success, more advanced networks have emerged including DIQaM [4], RankIQA [28], FPR [7], and GraphIQA [41]. However, limited data scales and learning strategies still restrict their generalization across different distributions. To account for this, some works leverage general knowledge learned from other tasks for IQA, such as image classification [57], image restoration [25], and image-language correspondence [38, 45]. Other attempts have been dedicated to the new learning paradigms, including meat learning [66], continual learning [65], curriculum learning [46], and domain adaptation/generalization [5, 6]. Despite the impressive progress, enhancing generalization in NR-IQA remains a significant challenge.
LMM-based IQA. Recently, high-performance LMMs such as ChatGPT-4v [37], LLaVA [26, 27], and mPLUG-Owl3 [56] have exhibited strong zero-shot transfer capabilities across various vision tasks, suggesting their potential applicability in the IQA task. Early efforts, such as Q-Bench [49] and DepictQA [58], established comprehensive benchmarks to evaluate LMMs in low-level visual perception. In [67], Zhu et al. examined the IQA performance of LMMs under the two-alternative forced choice (2AFC) prompting. Instead of score regression, Q-Align [50] enabled LMMs to rate images at text-defined rating levels, excelling in both quality and aesthetic assessment tasks. By utilizing pre-trained LMMs and extensive datasets, recent studies have continued to improve IQA outcomes [51, 52, 68]. Although fine-tuning LMMs for IQA yields strong performance, it also poses the risk of catastrophic forgetting [30] problems, leading to the degraded generalization capability on other tasks. In light of these limitations, training-free approaches offer a more promising alternative to fully harness the extensive knowledge in LMMs for IQA. However, this strategy remains relatively underexplored.
3 Methodology
3.1 Preliminary
@@
3.2 Framework of Perception Bias Mitigation
@@ 3.2.1 Bias Exposure
where $M(u,v)$ is the spatter mask indicating regions that are affected or unaffected, and $C(u,v)$ represents the specific color distribution adapted for different splatter types. Herein, we use the implementation in [18] to generate $M(u,v)$ and $C(u,v)$.
@@
where the $x^{F}$ is the fog pattern generated by a diamond-square algorithm [10, 18] based on $x$. $k$ is the hyper-parameter of distortion severity. As shown in Fig. 3, the four types of distortion degrade the image quality significantly while the semantics are still recognizable, leading to an expected bia exposure when they are not deemed poor quality by the LMM.
3.2.2 Bias Mitigation
@@ 4 Experiments
Figure 4: Visualization of image quality prediction results. In each subfigure, the top-left label shows numbers in green, blue, and red, representing the MOS, the LMM prediction result with the prompt in Q-Bench, and Our result, respectively.
4.1 Experimental Settings
@@
Datasets. We evaluate our method on six publicly available datasets: LIVE Challenge [14], KonIQ-10k [19], AGIQA-3k [23], KADID-10k [24], SPAQ [9] and SIQAD [55]. The KonIQ-10k, SPAQ, and LIVE Challenge datasets are in-the-wild image collections, featuring authentic distortions. KonIQ-10k and SPAQ datasets each contain over 10,000 images and the SPAQ dataset is specifically designed to assess the quality of images captured by smartphones. The KADID-10k dataset comprises 10,125 images with various systematic distortions. The AGIQA-3k dataset includes 2,900 images focused on AI-generated image quality assessment. The SIQAD dataset is constructed for screen content image (SCI) quality assessment and a total of 980 distorted SCIs are involved.
@@
Comparison Methods. We denote our method as “Q-Debias” and compare its performance against both training-free (opinion-unaware) and training-based methods across multiple datasets. The training-free methods include BLINDS-II [35], QAC [53], BRISQUE [33], NIQE [34], ILNIQE [62], NPQI [29], ContentSep [2], CLIP-IQA [45], MDFS [36], and Q-Bench [49]. In particular, Q-Bench refers to using the same LMM (mPLUG-Owl3) with our method, while applying the query prompt from Q-Bench. For the training-based methods, we adopt models trained on the large-scale KonIQ-10k dataset for comparison, including ARNIQA [1], TReS [15], and MUSIQ [22]. We list their performance in cross-dataset settings to verify their generalization capability. The Pearson Linear Correlation Coefficient (PLCC) and the Spearman Rank-Order Correlation Coefficient (SRCC) are used as metrics to assess the linearity and monotonicity of our quality predictions.
@@
Implementation Details. We select mPLUG-Owl3 as our LMM due to its superior performance on image processing tasks. We set $z_{i}\in\{1.00,1.01,1.02,\dots,1.10\}$, $n=11$ in Eqn. (3), and $c=2.0$ in Eqn. (6). We set $k=2.5$ in Eqn. (7). Our code is implemented in PyTorch and runs on an A6000 GPU.
@@
Table 1: Performance comparison of our method, Q-Debias, against both training-free and training-based IQA models. The Red percentage indicates the improvement of our method over Q-Bench. ‘Sp’ indicates results predicted when only the spatter-noise-degraded image is utilized as a condition under our enhancement framework. The best two results are highlighted in boldface.
@@
Methods Training-free? LIVE Challenge KonIQ-10k AGIQA-3k KADID-10k SPAQ SIQAD
(Training Set) SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$
BLIINDS-II [35] 0.090 0.107 0.585 0.598 0.454 0.510 0.224 0.313 0.317 0.326 0.681 0.726
QAC [53] 0.226 0.284 0.340 0.291 - - 0.239 0.309 0.440 0.450 0.300 0.375
BRISQUE [33] 0.561 0.598 0.705 0.707 0.493 0.533 0.330 0.370 0.484 0.481 0.157 0.215
NIQE [34] 0.463 0.491 0.551 0.488 0.528 0.520 0.379 0.389 0.703 0.671 0.369 0.341
ILNIQE [62] 0.439 0.503 0.505 0.496 0.594 0.623 0.540 0.534 0.696 0.637 0.375 0.385
NPQI [29] 0.475 0.490 0.613 0.614 0.658 0.714 0.391 0.340 0.600 0.616 - -
ContentSep [2] 0.506 0.513 0.640 0.627 - - 0.506 0.357 0.708 0.665 - -
CLIP-IQA [45] 0.612 0.594 0.695 0.727 0.658 0.714 0.500 0.520 0.738 0.735 0.596 0.657
MDFS [36] 0.482 0.536 0.733 0.712 0.672 0.676 0.598 0.594 0.741 0.718 - -
ARNIQA [1] ✗ (KonIQ-10k) 0.670 0.715 - - 0.621 0.694 0.725 0.717 0.576 0.577 0.558 0.591
TReS [15] ✗ (KonIQ-10k) 0.771 0.805 - - 0.652 0.737 0.468 0.492 0.418 0.417 0.572 0.625
MUSIQ [22] ✗ (KonIQ-10k) 0.788 0.824 - - 0.630 0.722 0.556 0.575 0.726 0.738 0.554 0.598
Q-Bench *  [49] 0.721 0.677 0.672 0.573 0.596 0.469 0.315 0.267 0.767 0.650 0.333 0.402
Q-Debias (Ours) 0.794 0.790 0.838 0.863 0.717 0.753 0.700 0.713 0.867 0.826 0.324 0.341
Improvement over Q-bench $\uparrow$ 10.1% $\uparrow$ 16.7% $\uparrow$ 24.7% $\uparrow$ 50.6% $\uparrow$ 20.3% $\uparrow$ 60.6% $\uparrow$ 122.2% $\uparrow$ 167.0% $\uparrow$ 13.0% $\uparrow$ 27.1% Sp: 0.385 $\uparrow$ 18.8% Sp: 0.427 $\uparrow$ 25.2%
@@
Vanilla prompt Our debias scheme Vanilla prompt Our debias scheme Vanilla prompt Our debias scheme
[…]
BakLLaVA [26] 0.090 0.108 0.263 ($\uparrow$ 192.0%) 0.265 ($\uparrow$ 145.0%) 0.480 0.321 0.460 0.482 ($\uparrow$ 50.2%) 0.285 0.215 0.362 ($\uparrow$ 27.0%) 0.374 ($\uparrow$ 74.0%)
Qwen-VL [3] 0.470 0.546 0.504 ($\uparrow$ 7.20%) 0.501 0.504 0.532 0.615 ($\uparrow$ 22.0%) 0.623 ($\uparrow$ 17.1%) 0.487 0.539 0.560 ($\uparrow$ 15.0%) 0.562 ($\uparrow$ 4.27%)
Table 3: Ablation study of different types of conditional images. The best two results are highlighted in boldface.
Exp. ID Setting LIVE Challenge SIQAD AGIQA-3k
[…]
1 0.721 0.677 0.333 0.402 0.596 0.469
2 $\checkmark$ $\checkmark$ 0.644 0.489 0.197 0.168 0.633 0.539
3 $\checkmark$ $\checkmark$ 0.762 0.694 0.385 0.427 0.713 0.704
4 $\checkmark$ $\checkmark$ 0.726 0.615 0.354 0.325 0.685 0.644
5 $\checkmark$ $\checkmark$ 0.784 0.790 0.265 0.258 0.720 0.735
6 $\checkmark$ $\checkmark$ $\checkmark$ 0.793 0.790 0.335 0.294 0.709 0.712
7 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.793 0.773 0.349 0.308 0.714 0.702
8 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.714 0.620 0.114 0.128 0.707 0.613
Q-Debias $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.794 0.790 0.326 0.341 0.717 0.753
4.2 Comparison with NR-IQA Models
Prediction Accuracy. As shown in Table 1, compared with existing training-free methods, our model Q-Debias consistently achieves the best performance across the first five IQA datasets. In particular, most hand-crafted feature-based models (e.g., NIQE) experience a significant performance drop on the KADID-10k dataset due to the diverse distortion types involved. However, our method still outperforms these models by a large margin, demonstrating its high effectiveness. Compared to the vanilla prompt used in Q-Bench, our method—while utilizing the same base LMM (mPLUG-Owl3)—achieves performance gains across all five datasets. Notably, both Q-Bench and our method do not achieve promising results on the SIQAD dataset. The reason may lie in that LMMs are primarily trained on natural images, while the SCIs in the SIQAD dataset have been less extensively explored. From our experiments, we observe that not all distortion types for conditional image generation are effective for SIQAD. However, the introduction of single spatter noise demonstrates a positive effect on the SIQAD dataset, suggesting that our conditional prompt strategy can still be effective when specific distortions are carefully explored.

### v2

@@ Abstract
Despite the impressive performance of large multimodal models (LMMs) in high-level visual tasks, their capacity for image quality assessment (IQA) remains limited. One main reason is that LMMs are primarily trained for high-level tasks (e.g., image captioning), emphasizing unified image semantics extraction under varied quality. Such semantic-aware yet quality-insensitive perception bias inevitably leads to a heavy reliance on image semantics when those LMMs are forced for quality rating. In this paper, instead of retraining or tuning an LMM costly, we propose a training-free debiasing framework, in which the image quality prediction is rectified by mitigating the bias caused by image semantics. Specifically, we first explore several semantic-preserving distortions that can significantly degrade image quality while maintaining identifiable semantics. By applying these specific distortions to the query/test images, we ensure that the degraded images are recognized as poor quality while their semantics mainly remain. During quality inference, both a query image and its corresponding degraded version are fed to the LMM along with a prompt indicating that the query image quality should be inferred under the condition that the degraded one is deemed poor quality. This prior condition effectively aligns the LMM’s quality perception, as all degraded images are consistently rated as poor quality, regardless of their semantic variance. Finally, the quality scores of the query image inferred under different prior conditions (degraded versions) are aggregated using a conditional probability model. Extensive experiments on various IQA datasets show that our debiasing framework could consistently enhance the LMM performance.
Codehttps://barrypan12138.github.io/Q-Debias/
Introduction
@@
Figure 1: Illustration of perception bias in Large Multimodal Model (LMM) during quality assessment. Image quality ratings from the LMM (mPLUG-Owl3 (Ye et al. 2024)) were obtained using the Q-Bench testing framework (Wu et al. 2023a). The LMM consistently assigns higher quality ratings to images in the second row compared to the first, despite both sets exhibiting similar quality distributions as measured by Mean Opinion Scores (MOSs). This discrepancy suggests that the LMM relies more on image semantics than on low-level image clues for quality assessment.
@@
No-Reference Image Quality Assessment (NR-IQA) models aim to measure image quality in alignment with human perception without any reference, playing a fundamental role across various computer vision tasks (Wang and Bovik 2006). In the past decades, significant progress has been witnessed in NR-IQA, including traditional hand-crafted feature-based models (Chen et al. 2025) and deep learning-based models (Chen et al. 2021). Despite the advancement, current models still grapple with limited generalization capabilities on unseen scenes and distortions due to the significant distribution shifts between training and test sets. Recently, the emergence of Large Multimodal Models (LMMs) (OpenAI 2023) has demonstrated impressive generalization abilities across various vision-language tasks, such as classification (Wang et al. 2023), image captioning (Yu et al. 2023), and visual question answering (Shang et al. 2024). However, the focus on high-level visual tasks usually limits their effectiveness in low-level tasks, such as IQA (Wu et al. 2023a; Sun et al. 2024). Although pioneering researchers have attempted to fine-tune or retrain LMMs for improved IQA accuracy (Wu et al. 2023b; Chen et al. 2024; Zhu et al. 2024), the laborious dataset construction and costly model training usually render this approach inefficient. Moreover, tuning LMMs specifically for IQA also introduces the risk of catastrophic forgetting (Luo et al. 2023), compromising the retention of general knowledge and ultimately degrading the models’ capability on other tasks.
@@
A powerful strategy to both retain the LMM’s strengths across tasks and enhance its IQA performance is to unlock its vast general knowledge through well-crafted prompts, encouraging the LMM to respond accurately to quality rating requests. However, despite this potential, LMMs, driven by training objectives that emphasize semantic extraction over quality, usually default to interpreting image quality heavily relying on image semantics. As shown in Fig. 1, two sets of distorted images with different semantics are presented to an LMM for quality rating. The results indicate that the LMM consistently prefers the quality of images with the second set of semantics, despite both sets having similar quality (Mean Opinion Score (MOS)) distributions. The case reveals that the LMM intrinsically bases its quality rating on image semantics rather than on quality-related clues (e.g., blur). Motivated by this observation, we introduce an innovative, training-free approach to mitigate the perception bias inherent in LMMs. Our enhancement strategy consists of two main steps: 1) bias exposure, and 2) bias mitigation.
@@
In the bias exposure step , we assume the bias exists consistently across images sharing the same semantics. Based on this assumption, we can expose the perception bias of a query/test image by measuring how much the LMM resists labeling it as high quality when its quality is severely degraded but the semantics are preserved. To achieve this, we explore several specific distortions that drastically corrupt the image quality while preserving its semantics to some extent. For a query image, we then impose these distortions on the query image and obtain its degraded versions. Herein, those degraded images should be deemed as poor quality. However, the LMM may not always agree with the fact and the disagreement leads to the bias exposure.
In the bias mitigation step, we address the exposed bias using an instructive conditional prompt. Specifically, during inference, we provide both the query image and its degraded version to the LMM, along with a prompt indicating that the quality of the query image should be assessed under the condition that the degraded counterpart is rated as poor quality. By forcing the LMM to rate the quality of the degraded images appropriately, the bias mitigation in turn refines the LMM’s quality prediction for the query image. Finally, the quality predictions under different distortions are aggregated through a conditional probability model, further improving the prediction accuracy. Before delving into detail, we highlight our main contributions as follows:
@@
Preliminary
@@
Framework of Perception Bias Mitigation
@@
where $M(u,v)$ is the spatter mask indicating regions that are affected or unaffected, and $C(u,v)$ represents the specific color distribution adapted for different splatter types. Herein, we use the implementation in (Hendrycks and Dietterich 2019) to generate $M(u,v)$ and $C(u,v)$.
@@
where the $x^{F}$ is the fog pattern generated by a diamond-square algorithm (Fournier, Fussell, and Carpenter 1982; Hendrycks and Dietterich 2019) based on $x$. $k$ is the hyper-parameter of distortion severity. As shown in Fig. 3, the four types of distortion degrade the image quality significantly while the semantics are still recognizable, leading to an expected bia exposure when they are not deemed poor quality by the LMM.
Bias Mitigation
@@
Figure 4: Visualization of image quality prediction results. In each subfigure, the top-left label shows numbers in green, blue and red, representing the MOS, the LMM prediction result with the prompt in Q-Bench and our result, respectively.
Experimental Settings
@@
Datasets. We evaluate our method on five publicly available datasets: LIVE Challenge (Ghadiyaram and Bovik 2015), KonIQ-10k (Hosu et al. 2020), AGIQA-3k (Li et al. 2023), KADID-10k (Lin, Hosu, and Saupe 2019), and SPAQ (Fang et al. 2020). The KonIQ-10k, SPAQ, and LIVE Challenge datasets are in-the-wild image collections, featuring authentic distortions. KonIQ-10k and SPAQ datasets each contain over 10,000 images and the SPAQ dataset is specifically designed to assess the quality of images captured by smartphones. The KADID-10k dataset comprises 10,125 images with various systematic distortions. The AGIQA-3k dataset includes 2,900 images focused on AI-generated image quality assessment.
@@
Comparison Methods. We denote our method as “Q-Debias” and compare its performance against both training-free (opinion-unaware) and training-based methods across multiple datasets. The training-free methods include BLINDS-II (Moorthy and Bovik 2010), BRISQUE (Mittal, Moorthy, and Bovik 2012), NIQE (Mittal, Soundararajan, and Bovik 2012), NPQI (Liu et al. 2020), ContentSep (Babu, Kannan, and Soundararajan 2023), CLIP-IQA (Wang, Chan, and Loy 2023), MDFS (Ni et al. 2024), and Q-Bench (Wu et al. 2023a). In particular, Q-Bench refers to using the same LMM (mPLUG-Owl3) with our method, while applying the query prompt from Q-Bench. For the training-based methods, we adopt models trained on the large-scale KonIQ-10k dataset for comparison, including ARNIQA (Agnolucci et al. 2024), TReS (Golestaneh, Dadsetan, and Kitani 2022), and MUSIQ (Ke et al. 2021). We list their performance in cross-dataset settings to verify their generalization capability. The Pearson Linear Correlation Coefficient (PLCC) and the Spearman Rank-Order Correlation Coefficient (SRCC) are used as metrics to assess the linearity and monotonicity of our quality predictions.
@@
Implementation Details. We select mPLUG-Owl3 as our LMM due to its superior performance on image processing tasks. We set $z_{i}\in\{1.00,1.01,1.02,\dots,1.10\}$, $n=11$ in Eqn. (3), and $c=2.0$ in Eqn. (6). We set $k=2.5$ in Eqn. (7).
@@
Table 1: Performance comparison of our method, Q-Debias, against both training-free and training-based IQA models. The percentage indicates the improvement of our method over Q-Bench. The best two results are highlighted in boldface.
@@
Methods Training-free? LIVE Challenge KonIQ-10k AGIQA-3k KADID-10k SPAQ
(Training set) SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑
BLIINDS-II (Moorthy and Bovik 2010) 0.090 0.107 0.585 0.598 0.454 0.510 0.224 0.313 0.317 0.326
QAC (Xue, Zhang, and Mou 2013) 0.226 0.284 0.340 0.291 - - 0.239 0.309 0.440 0.450
BRISQUE (Mittal, Moorthy, and Bovik 2012) 0.561 0.598 0.705 0.707 0.493 0.533 0.330 0.370 0.484 0.481
NIQE (Mittal, Soundararajan, and Bovik 2012) 0.463 0.491 0.551 0.488 0.528 0.520 0.379 0.389 0.703 0.671
ILNIQE (Zhang, Zhang, and Bovik 2015) 0.439 0.503 0.505 0.496 0.594 0.623 0.540 0.534 0.696 0.637
NPQI (Liu et al. 2020) 0.475 0.490 0.613 0.614 0.658 0.714 0.391 0.340 0.600 0.616
ContentSep (Babu, Kannan, and Soundararajan 2023) 0.506 0.513 0.640 0.627 - - 0.506 0.357 0.708 0.665
CLIP-IQA (Wang, Chan, and Loy 2023) 0.612 0.594 0.695 0.727 0.658 0.714 0.500 0.520 0.738 0.735
MDFS (Ni et al. 2024) 0.482 0.536 0.733 0.712 0.672 0.676 0.598 0.594 0.741 0.718
ARNIQA (Agnolucci et al. 2024) ✗ (KonIQ-10k) 0.670 0.715 - - 0.621 0.694 0.725 0.717 0.576 0.577
TReS (Golestaneh, Dadsetan, and Kitani 2022) ✗ (KonIQ-10k) 0.771 0.805 - - 0.652 0.737 0.468 0.492 0.418 0.417
MUSIQ (Ke et al. 2021) ✗ (KonIQ-10k) 0.788 0.824 - - 0.630 0.722 0.556 0.575 0.726 0.738
Q-Bench *  (Wu et al. 2023a) 0.721 0.677 0.672 0.573 0.596 0.469 0.315 0.267 0.767 0.650
Q-Debias (Ours) 0.794 0.818 0.838 0.863 0.717 0.753 0.700 0.753 0.867 0.826
Improvement over Q-bench $\uparrow$ 10.1% $\uparrow$ 20.8% $\uparrow$ 24.7% $\uparrow$ 50.6% $\uparrow$ 20.3% $\uparrow$ 60.6% $\uparrow$ 122.2% $\uparrow$ 167.0% $\uparrow$ 13.0% $\uparrow$ 27.1%
@@
Vanilla prompt Our debias prompt Vanilla prompt Our debias prompt Vanilla prompt Our debias prompt
SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑
BakLLaVA (Liu et al. 2024) 0.090 0.108 0.263 ( $\uparrow$ 192.0%) 0.265 ( $\uparrow$ 145.0%) 0.480 0.321 0.460 0.482 ( $\uparrow$ 50.2%) 0.285 0.215 0.362 ( $\uparrow$ 27.0%) 0.374 ( $\uparrow$ 74.0%)
Qwen-VL (Bai et al. 2023) 0.470 0.546 0.504 ( $\uparrow$ 7.20%) 0.501 0.504 0.532 0.615 ( $\uparrow$ 22.0%) 0.623 ( $\uparrow$ 17.1%) 0.487 0.539 0.560 ( $\uparrow$ 15.0%) 0.562 ( $\uparrow$ 4.27%)
LLaVA-OneVision (Li et al. 2024a) 0.379 0.654 0.631 ( $\uparrow$ 66.5%) 0.649 0.581 0.781 0.707 ( $\uparrow$ 8.94%) 0.806 ( $\uparrow$ 14.0%) 0.480 0.718 0.669 ( $\uparrow$ 39.4%) 0.728( $\uparrow$ 1.39%)
LLaVA-Interleave (Li et al. 2024b) 0.221 0.337 0.454 ( $\uparrow$ 100.5%) 0.543 ( $\uparrow$ 61.1%) 0.223 0.315 0.464 ( $\uparrow$ 108.1%) 0.560 ( $\uparrow$ 77.8%) 0.222 0.326 0.459 ( $\uparrow$ 106.8%) 0.552( $\uparrow$ 69.3%)
DeepSeek-VL2 (Wu et al. 2024) 0.800 0.851 0.822 ( $\uparrow$ 2.75%) 0.860 ( $\uparrow$ 1.06%) 0.606 0.655 0.759 ( $\uparrow$ 25.2%) 0.778 ( $\uparrow$ 17.0%) 0.703 0.753 0.791 ( $\uparrow$ 12.5%) 0.819( $\uparrow$ 8.76%)
Table 3: Ablation study of different types of conditional images. The best results in each setting are highlighted in boldface.
Exp. ID Single Distortion
Semantic
Consistency
LIVE Challenge AGIQA-3k Average
SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑
1 Blur Zoom $\checkmark$ 0.644 0.673 0.633 0.658 0.639 0.666
2 Motion $\checkmark$ 0.497 0.515 0.552 0.546 0.525 0.531
3 Gaussian $\checkmark$ 0.617 0.497 0.636 0.648 0.627 0.573
4 Noise Gaussian $\checkmark$ 0.677 0.646 0.686 0.720 0.682 0.683
5 Spatter $\checkmark$ 0.762 0.799 0.713 0.768 0.738 0.784
6 Bad weather Snow $\checkmark$ 0.713 0.761 0.686 0.640 0.700 0.701
7 Frost $\checkmark$ 0.632 0.705 0.633 0.573 0.633 0.639
8 Fog $\checkmark$ 0.729 0.763 0.689 0.702 0.709 0.733
9 Brightness $\checkmark$ 0.613 0.673 0.620 0.668 0.617 0.671
10 Saturation $\checkmark$ 0.784 0.790 0.720 0.735 0.752 0.763
Multiple Distortions
Exp. ID Zoom Spatter Fog Saturation Semantic SRCC↑ PLCC↑ SRCC↑ PLCC↑ SRCC↑ PLCC↑
11 0.721 0.677 0.596 0.469 0.659 0.573
12 $\checkmark$ $\checkmark$ $\checkmark$ 0.793 0.790 0.709 0.712 0.751 0.763
13 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.793 0.773 0.714 0.702 0.753 0.738
14 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.493 0.472 0.518 0.508 0.506 0.490
Q-Debias $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.794 0.818 0.717 0.753 0.756 0.786
Comparison with NR-IQA Models
Prediction Accuracy. As shown in Table 1, compared with existing training-free methods, our model Q-Debias consistently achieves the best performance across the first five IQA datasets. In particular, most hand-crafted feature-based models (e.g., NIQE) experience a significant performance drop on the KADID-10k dataset due to the diverse distortion types involved. However, our method still outperforms these models by a large margin, demonstrating its high effectiveness. Compared to the vanilla prompt used in Q-Bench, our method while utilizing the same base LMM (mPLUG-Owl3), achieves performance gains across all five datasets.